OpenStudy (anonymous):

How can I figure out this question? Use linear approximation i.e. tangent line to approximate square root of 49.1 as follows, Let f(x0 = square root of x and the equation of the tangent line to f(x) can be written in the form x=49 in the form y=mx+b. where m is: and b is: Using this we find our approximation for square root of 49.1 is...

OpenStudy (anonymous):

The linear approximation to a function \(f(x)\) at a point \(c\) is \[f(x)\approx f'(c)(x-c)+f(c)\] To approximate \(\sqrt{49.1}\), you would set \(x=49.1\) and \(c\) to some value whose square root is easy to find and close to the given \(x\). In this case, \(c=49\). \[f(x)=\sqrt{x}~~\Rightarrow~~f'(x)=\frac{1}{2\sqrt x}\\ f(49.1)\approx f'(49)(49.1-49)+\sqrt{49}\\ ~~~~~~~~~~~~=\frac{1}{2\sqrt{49}}(0.1)+7\\ ~~~~~~~~~~~~=\cdots\]